[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jun 27, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Jun 27 13:00:46 UTC 2016


Here is today's summary of NHC computer operations:
--- mello

----------

TAFB:
no new issues.
HSU:
HSU reported the hotline did not ringing on our end late last week. The  
hurricane hotline had no issues with the medium-range call-- it rang  
internally in a normal fashion over the weekend.
Other items:
Salime fixed NHC's problem with the Local ingest of GOES-E. It appears the  
roof work resulted in a dirty satellite fiber optic cable. TSB will switch  
the GOES-E data feed back on Today or Tuesday.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160627/63da7f25/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list